OpenStudy (anonymous):

simplify (sqrt -3)^3 PLEASE HELP :(

OpenStudy (mertsj):

\[(\sqrt{-3})^3=(\sqrt{-3})(\sqrt{-3})(\sqrt{-3})=(i \sqrt{3})(i \sqrt{3})(i \sqrt{3})=i^3(3\sqrt{3})=-3i \sqrt{3}\]